How do you find x and y of this rectangle when sides:
AB = -2x
DC = 7y + 1
AD = -x
BC = 2y + 2
AB = DC
AD = BC
I know you use substitution, I'm just not sure how to... Thank you so much in advance!

Answer by ankor@dixie-net.com(22740)

How do you find x and y of this rectangle when sides:
AB = -2x
DC = 7y + 1
AD = -x
BC = 2y + 2
AB = DC
AD = BC
:
Draw out a rectangle ABCD, label each side as given:
Put the equations in the standard form
AB=DC, therefore:
-2x = 7y+1
-2x - 7y = 1
2x + 7y = -1; multiplied by -1
and
AD=BC, therefore
-x = 2y+2
-x - 2y = 2
x + 2y = -2
:
Use these two equation for elimination
2x + 7y = -1
x + 2y = -2
multiply the 2nd equation by 2
2x + 7y = -1
2x + 4y = -4
---------------Subtracting eliminates x, find y
3y = 3
y = 1
:
You should be able to find x, check to make sure opposite sides are equal